# DyPE: Dynamic Position Extrapolation for Ultra High Resolution Diffusion

### v1

@@ 1 Introduction
Diffusion Transformers (DiTs) (Peebles & Xie, 2022) have recently emerged as a powerful class of generative models, combining the stable training dynamics of diffusion (Ho et al., 2020; Song et al., 2020) with the expressiveness and scalability of transformers (Vaswani et al., 2017; Kaplan et al., 2020; Hoffmann et al., 2022). While this architecture fueled progress across large-scale vision (Dosovitskiy et al., 2021), training these models to ultra-high resolutions (e.g., $4096^{2}$ and beyond) remains a formidable challenge: the quadratic complexity of self-attention in the number of image tokens drives up memory and compute costs, making direct training infeasible.
This limitation is analogous to the long-context challenge in large language models (LLMs), where transformers are trained with a fixed context horizon, but are expected to perform on much longer sequences during inference. The positional encoding (PE) mechanism is central to this generalization, as it dictates how transformers align and extrapolate positional relations across unseen ranges. Rotary Positional Embeddings (RoPE) (Su et al., 2021) are widely adopted but degrade when extrapolated beyond the training range. This has motivated inference-time adaptations such as position interpolation (PI) (Chen et al., 2023b), NTK-aware rescaling (Peng et al., 2023a), and YaRN (Peng et al., 2023b), which adjust the frequency spectrum to better preserve long-range dependencies.
In image generation, these LLM-derived schemes were adapted to accommodate aspect-ratio changes and moderate increases in resolution (Lu et al., 2024; YourTeam, 2025). However, these static approaches do not account for the distinctive spectral progression of the diffusion process, where low-frequency structures are generated in the first sampling steps, while high-frequency details are resolved later (Rissanen et al., 2023; Hoogeboom et al., 2023; Chen et al., 2023c). As shown in Zhuo et al. (2024), aligning with these dynamics can facilitate better resolution extrapolation. These observations naturally lead to our guiding question: How should positional embeddings be dynamically adjusted to reflect the spectral progression of the diffusion process?
@@
This training-free strategy greatly improves generalization, allowing a pre-trained FLUX model (Lee et al., 2025) to generate images at ultra-high resolutions (exceeding 1616M pixels), as shown in Fig. 1. We evaluate DyPE using quantitative metrics for image quality and prompt fidelity, alongside qualitative and human evaluations. The results show that DyPE achieves consistent improvements in ultra-high-resolution synthesis across multiple benchmarks and resolutions, all without retraining or additional sampling costs.
@@ 2.2 Rotary Positional Embeddings and Position Extrapolation
The transformer block, which is the basis of DiT, is permutation equivariant. Thus, a positional encoding mechanism is necessary to properly model the strong spatial dependencies in natural images (LeCun & Bengio, 1998). Early solutions use fixed sinusoidal positional embedding (Vaswani et al., 2017; Dosovitskiy et al., 2021), learned absolute embeddings (Devlin et al., 2019; Radford et al., 2019), or relative positional embeddings (Press et al., 2021). More recently, the Rotary Positional Embeddings (RoPE) (Su et al., 2021) emerged as a more effective alternative which provides the relative positions in the query–key interactions.
@@
with corresponding wavelength $\lambda_{d}=2\pi/\theta_{d}$, where $\theta_{\text{base}}$ is a model hyper-parameter.
We note that in case of 22D images RoPE is applied axially: half of the hidden vector is rotated horizontally, and the other half vertically. Thus this axial decomposition enables RoPE to encode relative offsets along each axis independently, considering the spatial structure of images (Heo et al., 2024).
As discussed above, training DiT models at high resolutions incurs substantial memory and compute cost. Applying a model at higher resolutions than it was trained on, suffers from degraded performance as illustrated in Fig. 1. This shortcoming spurred the development of inference-time positional encoding adaptations for a better generalization. Before we survey these approaches, let us establish useful notations from Peng et al. (2023b).
@@
This mapping resamples the waves $\cos(m\theta_{d}),\sin(m\theta_{d})$ at a finer rate in the larger context grid $L^{\prime}$, and while it correctly reproduces the lower end of the spectrum, it fails to reach the new grid’s higher frequency band. While large scale content is properly synthesizes in this approach, the missing high-frequencies manifest as blurriness and lack of fine detail, as discussed in Appendix A.
NTK-Aware Interpolation. To address this problem, the Neural Tangent Kernel (NTK-aware) interpolation (Peng et al., 2023a; b) applies different scaling to the low and high frequencies, by:
@@
YaRN. Yet another RoPE extensioN, or YaRN (Peng et al., 2023b) extends the latter in two ways. The first is the NTK-by-parts interpolation, which splits the spectrum to three bands, where different mappings are applied, namely:
@@ 3.1 Evolution of Frequency Modes in the Diffusion Process
where $\hat{(\cdot)}$ denotes the Fourier transformed signals. The i.i.d noise vectors $\epsilon$ have a white (constant) Power Spectrum Density (PSD), and the data of natural images, $x_{t}$, is known to have a well-characterized PSD with a power-law decay of $\propto 1/f^{\omega}$ where $\omega\approx 2$ (van der Schaaf & van Hateren, 1996; Hyvrinen et al., 2009), as function of frequency $f$. These terms allow us to explicitly describe the time-dependent mean PSD in Eq. 10, given by
@@
This topic is briefly touched by Zhuo et al. (2024) as part of deriving a new DiT architecture. However, the opposite conclusions are drawn. We discuss this strategy in Appendix B.
@@ 4 Experiments
Figure 3: Zoom-in comparison at $4096^{2}$ of Dy-YaRN vs. YaRN. Three magnified regions highlight fine-detail differences. Additional example can be found in Fig. 13 in the Appendix.
@@
We first apply DyPE on top of FLUX (Lee et al., 2025), with evaluations on two established benchmarks, DrawBench (Saharia et al., 2022) and Aesthetic-4K (Zhang et al., 2025), including automatic metrics, human evaluation, and resolution-scaling analysis (Sec. 4.1). We then extend evaluation to class-conditional image synthesis on FiTv2 (Wang et al., 2024) (Sec. 4.2). We also include zoom-in studies to highlight improvements in preserving high-frequency details (Fig. 3). Furthermore, in Appendix D, we present an ablation study examining design choices, focusing on (i) scheduler variants for Dy-NTK-aware and (ii) timestep incorporation strategies for Dy-YaRN. Finally, additional results are provided in Appendix E, including further comparisons with baselines, panorama generation, and more visual examples. Implementation details are in Appendix C.
@@ 4.1 Ultra-High-Resolution Text-to-Image Generation
Table 1: High-resolution image generation on DrawBench and Aesthetic-4K, shown as two resolutions per row. Each row reports CLIPScore (CLIP), ImageReward (IR), Aesthetics (Aesth) for DrawBench, and CLIP, IR, Aesth, and FID for Aesthetic-4K. All methods are built on FLUX.
@@
Figure 4: Qualitative results at $4096^{2}$ resolution using two representative prompts from Aesthetic-4K. We compare NTK-aware, Dy-NTK-aware, YaRN, and Dy-YaRN.
@@
Baselines. We evaluate DyPE on top of the pre-trained FLUX (Lee et al., 2025), specifically the FLUX.1-Krea-dev version, whose effective generation resolution is $1024\times 1024$. As baselines, we use FLUX itself, and also, in test time only, apply on top of FLUX the positional embedding extrapolation methods NTK-aware and YaRN, adapted for vision tasks by applying them independently on the $x$ and $y$ axes. On top of these, we apply our DyPE, specifically Dy-NTK-aware, and Dy-YaRN.
@@
Benchmarks. As for benchmarks, we first consider DrawBench (Saharia et al., 2022), a set of 200 text prompts for evaluating text-to-image models across multiple criteria. Following Ma et al. (2025); Chachy et al. (2025), we measure: (i) text-image alignment using CLIP-Score (Hessel et al., 2022), a similarity metric between image and text embeddings based on CLIP (Radford et al., 2021), (ii) human preference alignment using ImageReward (Xu et al., 2023), a reward model trained on large-scale human feedback for generated images, and (iii) image aesthetics using Aesthetic-Score-Predictor (Schuhmann et al., 2022), a model trained to predict human aesthetic judgments. Additionally, to specifically assess fine-grained, ultra-high-resolution fidelity, we evaluate on Aesthetic-4K (Zhang et al., 2025). We use its 4K subset (Aesthetic-Eval@4096), which comprises 195 curated image–prompt pairs, and downsample them to match the target test resolutions for fair comparison. Following the official protocol, we report (i) CLIPScore, (ii) ImageReward, (iii) Aesthetics score, and (iv) FID (Heusel et al., 2017), which assesses the fidelity and diversity of generated images based on the distributional distance between real and generated features.
@@
Results. Quantitative results across different resolutions and aspect ratios are presented in Tab. 1, with Fig. 4 showing side-by-side comparisons on Aesthetic-4K. Additional qualitative results are provided in the Appendix for DrawBench (Fig. 11) and Aesthetic-4K (Fig. 12). As can be seen in Fig. 1, FLUX exhibited repeating artifacts in ultra-high-resolution revealing the periodicity of the sine waves in the larger context as further illustrated in Appendix. A. We also note that FLUX shows relatively stronger performance on landscape resolutions compared to portrait, likely reflecting a training bias, the gap on portrait settings widens once DyPE is applied, suggesting that our approach helps to unlock this limitation. Importantly, the advantage of DyPE becomes more pronounced as the generation resolution increases (e.g., to $3072^{2}$ and $4096^{2}$), underscoring the effectiveness of DyPE within diffusion transformers for ultra-high-resolution generation. Further visual results are presented in the supplementary.
@@
To complement the automatic metrics, we conduct a human study on a curated subset of 20 prompts from Aesthetic-4K, obtained by sampling every fourth entry to ensure uniform coverage. We consider 50 raters and present them with pairwise comparisons at $4096^{2}$ resolution, generated on FLUX. Each prompt yields two comparisons: (i) NTK-aware vs. Dy-NTK-aware, and (ii) YaRN vs. Dy-YaRN. For each pair, participants answer the following three questions: (i) Which image is more aligned with the given text prompt? (ii) Which image has better overall geometry and structure (coherent shapes, correct proportions, fewer distortions) and (iii) Which image has more aesthetic and realistic textures and fine details? Results, summarized in Tab. 2 shows that DyPE consistently achieves superior quality, with preference rates ranging from about 87% to nearly 90%.
@@
Figure 5: Resolution scaling analysis.

### v2

@@ Abstract
Machine Learning, ICML
The Hebrew University of Jerusalem
@@ 1 Introduction
Diffusion Transformers (DiTs) (Peebles and Xie, 2022) have recently emerged as a powerful class of generative models, combining the stable training dynamics of diffusion (Ho et al., 2020; Song et al., 2020) with the expressiveness and scalability of transformers (Vaswani et al., 2017; Kaplan et al., 2020; Hoffmann et al., 2022). While this architecture fueled progress across large-scale vision (Dosovitskiy et al., 2021), training these models to ultra-high resolutions (e.g., $4096^{2}$ and beyond) remains a formidable challenge: the quadratic complexity of self-attention in the number of image tokens drives up memory and compute costs, making direct training infeasible.
This limitation is analogous to the long-context challenge in large language models (LLMs), where transformers are trained with a fixed context horizon, but are expected to perform on much longer sequences during inference. The positional encoding (PE) mechanism is central to this generalization, as it dictates how transformers align and extrapolate positional relations across unseen ranges. Rotary Positional Embeddings (RoPE) (Su et al., 2021) are widely adopted but degrade when extrapolated beyond the training range. This has motivated inference-time adaptations such as position interpolation (PI) (Chen et al., 2023b), NTK-aware rescaling (Peng et al., 2023), and YaRN (Peng et al., 2023), which adjust the frequency spectrum to better preserve long-range dependencies.
In image generation, these LLM-derived schemes were adapted to accommodate aspect-ratio changes and moderate increases in resolution (Lu et al., 2024; Wang et al., 2024). However, these static approaches do not account for the distinctive spectral progression of the diffusion process, where low-frequency structures are generated in the first sampling steps, while high-frequency details are resolved later (Rissanen et al., 2023; Hoogeboom et al., 2023). As shown in Zhuo et al. (2024), aligning with these dynamics can facilitate better resolution extrapolation. These observations naturally lead to our guiding question: How should positional embeddings be dynamically adjusted to reflect the spectral progression of the diffusion process?
@@
This training-free strategy greatly improves generalization, allowing a pre-trained FLUX model (Labs, 2024) to generate images at ultra-high resolutions (exceeding 1616M pixels), as shown in Fig. 1. We evaluate DyPE using quantitative metrics for image quality and prompt fidelity, alongside qualitative and human evaluations. The results show that DyPE achieves consistent improvements in ultra-high-resolution synthesis across multiple benchmarks and resolutions, all without retraining or additional sampling costs.
@@ 2.2 Rotary Positional Embeddings and Position Extrapolation
The transformer block, which is the basis of DiT, is permutation equivariant. Thus, a positional encoding mechanism is necessary to properly model the strong spatial dependencies in natural images (LeCun and Bengio, 1998). Early solutions use fixed sinusoidal positional embedding (Vaswani et al., 2017; Dosovitskiy et al., 2021), learned absolute embeddings (Devlin et al., 2019; Radford et al., 2019), or relative positional embeddings (Press et al., 2021). More recently, the Rotary Positional Embeddings (RoPE) (Su et al., 2021) emerged as a more effective alternative which provides the relative positions in the query–key interactions.
@@
with corresponding wavelength $\lambda_{d}=2\pi/\theta_{d}$, where $\theta_{\text{base}}$ is a model hyper-parameter. We note that in case of 22D images RoPE is applied axially: half of the hidden vector is rotated horizontally, and the other half vertically. Thus this axial decomposition enables RoPE to encode relative offsets along each axis independently, considering the spatial structure of images (Heo et al., 2024).
As discussed above, training DiT models at high resolutions incurs substantial memory and compute cost. Applying a model at higher resolutions than it was trained on, suffers from degraded performance as illustrated in Fig. 1. This shortcoming spurred the development of inference-time positional encoding adaptations for a better generalization. Before we survey these approaches, let us establish useful notations from Peng et al. (2023).
@@
This mapping resamples the waves $\cos(m\theta_{d}),\sin(m\theta_{d})$ at a finer rate in the larger context grid $L^{\prime}$, and while it correctly reproduces the lower end of the spectrum, it fails to reach the new grid’s higher frequency band. While large scale content is properly synthesized in this approach, the missing high-frequencies manifest as blurriness and lack of fine detail, as discussed in Appendix A.
NTK-Aware Interpolation. To address this problem, the Neural Tangent Kernel (NTK-aware) interpolation (Peng et al., 2023) applies different scaling to the low and high frequencies, by:
@@
YaRN. Yet another RoPE extensioN, or YaRN (Peng et al., 2023) extends the latter in two ways. The first is the NTK-by-parts interpolation, which splits the spectrum to three bands, where different mappings are applied, namely:
@@ 3.1 Evolution of Frequency Modes in the Diffusion Process
where $\hat{(\cdot)}$ denotes the Fourier transformed signals. The i.i.d noise vectors $\epsilon$ have a white (constant) Power Spectrum Density (PSD), and the data of natural images, $x_{t}$, is known to have a well-characterized PSD with a power-law decay of $\propto 1/f^{\omega}$ where $\omega\approx 2$ (van der Schaaf and van Hateren, 1996; Hyvrinen et al., 2009), as function of frequency $f$. These terms allow us to explicitly describe the time-dependent mean PSD in Eq. 10, given by
@@
This topic is briefly touched by (Zhuo et al., 2024) as part of deriving a new DiT architecture. However, the opposite conclusions are drawn. We discuss this strategy in Appendix B.
@@ 4 Experiments
Figure 3: Zoom-in comparison at $4096^{2}$ of Dy-YaRN vs. YaRN. Additional example can be found in Fig. 17 in the Appendix.
@@
We first apply DyPE on top of FLUX (Labs, 2024), with evaluations on two established benchmarks, DrawBench (Saharia et al., 2022) and Aesthetic-4K (Zhang et al., 2025a), including automatic metrics, human evaluation, and resolution-scaling analysis (Sec. 4.1). We then extend evaluation to class-conditional image synthesis on FiTv2 (Wang et al., 2024) (Sec. 4.2). We also include zoom-in studies to highlight improvements in preserving high-frequency details (Fig. 3). Furthermore, in Appendix D, we present an ablation study examining design choices, focusing on (i) scheduler variants for Dy-NTK-aware and (ii) timestep incorporation strategies for Dy-YaRN. Finally, additional results are provided in Appendix E, covering additional DiT-based architectures (Qwen-Image (Wu et al., 2025)), high-resolution video generation (Wan et al., 2025), and high-resolution image editing tasks, panorama generation, and more visual examples. Implementation details are provided in Appendix C.
@@ 4.1 Ultra-High-Resolution Text-to-Image Generation
In Sec. 4.1.1 we evaluate DyPE against Position-Extapolation-based approaches and in Sec. 4.1.2 we evaluate DyPE against more general baselines as custom in previous works (Bu et al., 2025).
4.1.1 Comparison with Position-Extrapolation Baselines
Table 1: High-resolution image generation on DrawBench and Aesthetic-4K on multiple resolutions. Each row reports CLIPScore (CLIP), ImageReward (IR), Aesthetics (Aesth) for DrawBench, and CLIP, IR, Aesth, and FID for Aesthetic-4K. All methods are built on FLUX.
@@ 4.1.1 Comparison with Position-Extrapolation Baselines
Figure 4: Qualitative results at $4096^{2}$ resolution using two representative prompts from Aesthetic-4K.
@@
We evaluate DyPE on top of the pre-trained FLUX (Labs, 2024), specifically the FLUX.1-Krea-dev version, whose effective generation resolution is $1024\times 1024$. As primary baselines, we use FLUX itself and, in test time only, apply on top of FLUX the positional–embedding extrapolation methods NTK-aware and YaRN, adapted to vision by applying them independently on the $x$ and $y$ axes. We also compare with Time-Aware Scaled RoPE (TASR) (Zhuo et al., 2024), which interpolates from PI to NTK-aware scaling as denoising advances (discussed in Appendix B). On top of these, we evaluate our DyPE, including both Dy-NTK-aware and Dy-YaRN.
@@
Benchmarks. As for benchmarks, we first consider DrawBench (Saharia et al., 2022), a set of 200 text prompts for evaluating text-to-image models across multiple criteria. Following Ma et al. (2025); Chachy et al. (2025), we measure: (i) text-image alignment using CLIP-Score (Hessel et al., 2022), a similarity metric between image and text embeddings based on CLIP (Radford et al., 2021), (ii) human preference alignment using ImageReward (Xu et al., 2023), a reward model trained on large-scale human feedback for generated images, and (iii) image aesthetics using Aesthetic-Score-Predictor (Schuhmann et al., 2022), a model trained to predict human aesthetic judgments. Additionally, to specifically assess fine-grained, ultra-high-resolution fidelity, we evaluate on Aesthetic-4K (Zhang et al., 2025a). We use its 4K subset (Aesthetic-Eval@4096), which comprises 195 curated image–prompt pairs, and downsample them to match the target test resolutions for fair comparison. Following the official protocol, we report (i) CLIPScore, (ii) ImageReward, (iii) Aesthetics score, and (iv) FID (Heusel et al., 2017), which assesses the fidelity and diversity of generated images based on the distributional distance between real and generated features.
@@
Results. Quantitative results across different resolutions and aspect ratios are presented in Tab. 1, with Fig. 4 showing side-by-side comparisons on Aesthetic-4K. Additional qualitative results are provided in the Appendix for DrawBench (Fig. 15) and Aesthetic-4K (Fig. 16). As can be seen in Fig. 1, FLUX exhibits repeating artifacts at ultra-high resolutions, revealing the periodicity of the sinusoidal positional encoding when extrapolated to larger spatial contexts, as further illustrated in Appendix A. We also observe that FLUX performs relatively better on landscape resolutions than portrait, likely reflecting a training-set bias. Notably, once DyPE is applied, this gap widens in favor of our approach on portrait settings as well, indicating that DyPE helps mitigate this limitation. Importantly, the advantage of DyPE becomes increasingly pronounced as the generation resolution grows (e.g., up to $3072^{2}$ and $4096^{2}$), underscoring the effectiveness of our method for ultra-high-resolution synthesis in diffusion transformers. Further visual results are presented in the Appendix.
@@
To complement the automatic metrics, we conduct a human study on a curated subset of 20 prompts from Aesthetic-4K, obtained by sampling every fourth entry to ensure uniform coverage. We consider 50 raters and present them with pairwise comparisons at $4096^{2}$ resolution, generated on FLUX. Each prompt yields three comparisons: (i) NTK-aware vs. Dy-NTK-aware, (ii) Time-Aware Scaled RoPE (TASR) vs. Dy-NTK-aware, and (iii) YaRN vs. Dy-YaRN. For each pair, participants answer the following three questions: (i) Which image is more aligned with the given text prompt? (ii) Which image has better overall geometry and structure (coherent shapes, correct proportions, fewer distortions) and (iii) Which image has more aesthetic and realistic textures and fine details? Results, summarized in Tab. 2 shows that DyPE consistently achieves superior quality, with preference rates ranging from about 70.5% to 94.2%.
@@
Figure 5: Scaling analysis.
@@
4.1.2 Comparison with Other Approaches — General Baselines
Table 3: Evaluation at $2048^{2}$ and $4096^{2}$ resolutions on non-PE approaches. Baselines include U-Net-based methods (*) , tuning-based approaches (§), progressive refinement pipelines (‡), and diffusion combined with super-resolution (†).
$2048\times 2048$ $4096\times 4096$
Method FID↓ patch-FID↓ IS↑ patch-IS↑ FID↓ patch-FID↓ IS↑ patch-IS↑
DemoFusion* 205.59 199.00 10.03 10.49 205.86 195.69 10.93 7.92
FreCaS* 201.07 195.73 11.45 10.52 200.95 202.41 11.14 8.02
DiffuseHigh* 178.08 117.43 14.47 10.68 186.25 96.99 12.62 7.56
FreeScale* 199.87 126.45 9.89 10.55 259.24 191.23 10.66 7.74
[…]
Dy-YaRN (Ours) 142.74 96.34 14.76 11.95 186.00 78.33 13.96 10.13
In addition to PE approaches, we also compare with a broad set of non-PE methods spanning three families of high-resolution diffusion pipelines:
  1. (i)
    DiT-based methods - UltraPixel (Ren et al., 2024) (which relies on SD3 (Esser et al., 2024a)), Diffusion-4K (Zhang et al., 2025a) (which requires model fine-tuning), HiFlow (Bu et al., 2025), and I-Max (Du et al., 2024b) (both of which rely on multi-stage or progressive upscaling).
  2. (ii)
    U-Net–based methods - DemoFusion (Du et al., 2024a), FreCaS (Zhang et al., 2025b), DiffuseHigh (Kim et al., 2025), and FreeScale (Qiu et al., 2025).
  3. (iii)
    Diffusion + Super-Resolution - FLUX combined with BSRGAN (Zhang et al., 2021).
Benchmarks. Consistent with previous works (Bu et al., 2025), our evaluation focuses on patch-based fidelity and detail preservation. Specifically, we report: (i) FID (Heusel et al., 2017), (ii) patch-FID, (iii) Inception Score (IS) (Salimans et al., 2016), and (iv) patch-Inception Score (patch-IS). These complementary metrics isolate local structure and texture quality at ultra-high resolutions, allowing us to more precisely assess how well each method maintains fine-grained detail.
Results. We evaluate at resolutions $2048^{2}$ and $4096^{2}$ to specifically assess local detail preservation at extreme scales. As can be seen in Tab. 3, at $2048^{2}$ our Dy-YaRN variant achieves the best performance across all four metrics among all compared methods. At $4096^{2}$, Dy-YaRN attains the best patch-FID and patch-IS, while the Dy-YaRN+HiFlow combination yields the best global FID and IS. Notably, for every metric at both resolutions, at least one DyPE-based variant (either Dy-YaRN or Dy-YaRN+HiFlow) outperforms the baselines, highlighting the effectiveness of our approach. Additionally, in the Appendix, Fig. 14, we include qualitative comparison of our Dy-YaRN variant with representative baselines.